create a hypotesis based on the following

Several human physical traits are both distinctive and easily observed. Like tongue rolling, they are caused by the presence of either dominant or recessive alleles inherited from parents. You will survey people you know to obtain data on the presence of dominant and recessive alleles.

question

Based on the purpose and the background information, how do you think dominant traits will compare to recessive traits in your sample population?
Answer #1

A trait is a distinguishing characteristic of any individual. Whether a trait is dominant or recessive is determined by the allele that controls it. Each sexually reproducing organism carries two sets of chromosomes with each set contributed from each parent. Each set of chromosomes carries information for each and every trait of that organism and that information is encoded by genes.

Genes have a definite position on chromosomes which is called locus. Both sets of chromosomes may have different or same versions of these genes. However, only one gene is expressed out of these two. The one that is expressed over the other is termed as a dominant allele and the other is termed as a recessive allele.

Dominant traits are always expressed, but recessive traits are expressed only when both alleles for the trait are recessive. Dominant alleles are denoted by a capital letter (e.g., A), while recessive alleles are denoted by a small letter(e.g., a). When both alleles for a trait are same, the condition is known as homozygous condition (AA) and if they are different the condition is called heterozygous condition (Aa).

This can be further explained as below :

Dominant allele + dominant allele = dominant trait

A + A = AA

Dominant allele + recessive allele = dominant trait

A + a = Aa

Recessive allele + recessive allele = recessive trait

a + a = a

We can conclude that in any sample population, dominant traits will express over recessive traits. For example, if two people with a recessive trait marry and have children, all their children would have that recessive trait. On the other hand, if one has a dominant trait and is homozygous, their children would have that dominant trait. In heterozygous condition, they may have recessive or dominant trait depending on the presence of dominant or recessive allele.